# Limiting noise and outliers in the MICROSCOPE space test of the weak equivalence principle

### v3

@@ III Limiting noise, patch charges and fluctuating zero
Using (5), it obeys the relationship
@@
As recalled in Sec. II, the difference of acceleration biases between the TMs of SUEP is $\simeq 10^{-6}\,\rm ms^{-2}$. According to (8), a patch fluctuating potential $\delta V_{fpp}\simeq 0.3\,\mu V$ at the V3 signal frequency $\nu_{{}_{EP}}=3.11\,\rm mHz$ would produce a spurious acceleration of $4\times 10^{-14}\,\rm ms^{-2}$, as large as the average acceleration difference measured in all SUEP sessions at $\nu_{{}_{EP}}$ (ChhunCQG2022-5 , Fig. 10) and enough to change it by $100\%$ either way.
In GOCE a source of low frequency acceleration noise with $1/\sqrt{\nu}$ dependence below $0.01\,\rm Hz$ was found to be due to fluctuating potentials generated by charge patches on the test mass and its electrodesE2ESimulatorGOCE . GOCE was set up as a gravity gradiometer with $0.5\,\rm m$ separation of the TMs and a measurement bandwidth from $500\,\rm mHz$ to $0.1\,\rm Hz$. At $3\,\rm mHz$ the acceleration difference sensitivity was of $2\times 10^{-12}\,\rm ms^{-2}/\sqrt{Hz}$ along the ultra sensitive $Y,\,Z$ axes (more than one order of magnitude better than that of SUEP sensitive axis) and $4\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ along the least sensitive $X$ axis, with a contribution from patch fluctuating potentials of about $4\times 10^{-13}\,\rm ms^{-2}/\sqrt{Hz}$ and $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ respectively TouboulReview2016 .
@@
An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments, including those with laser tracked satellites, celestial bodies and cold atoms Blaser2001 ; Comandi2003 ; Nobili2008 ; Nobili2016 . Instead, in torsion-balance tests and the proposed GG space experiment the physical observable is a true null PerBraginskyPLA2018 .
IV Outliers, gaps and artificial data
@@
Since 2001 five space missions carrying ONERA’s accelerometers have been launched –CHAMP, GRACE, GOCE, GRACE Follow-On (GFO) and MICROSCOPE– with a total of seven satellites, and all but one have experienced such spikes (named “clanks” in CHAMP and GOCE, “twangs” in GRACE, “glitches” in MICROSCOPE). At the time of the design of GOCE, reports of such effects in CHAMP caused alarm, and countermeasures were adopted at design and test level, which were successful, and no spikes were seen. The succeeding missions discovered the problem anew.
Although as far as we know a comprehensive analysis across all the missions has not been carried out, there seems to be a consensus that the spikes are triggered by energy inputs from the Earth and the Sun causing micro-vibration, e.g. release of stress energy in the spacecraft materials such as the multi-layer insulation (MLI) GRACEglitchesTwangs1 ; GRACEglitchesTwangs2 . Alternative hypotheses point to small electric discharges in the spacecraft surfaces, or changes in the solar array currents. The GOCE countermeasures program was based on the mechanical hypothesis.
@@
The synodic frequency relative to the Sun is well approximated by the spin frequency relative to the fixed stars, the difference being the annual orbital frequency of the Earth around the Sun of $2\times 10^{-7}\,\rm rad/s$. MICROSCOPE spins around the $Y$ axis, perpendicular to the orbital plane, the spin frequency is opposite to the orbital frequency and is set at an odd multiple $N$ of half the orbital frequency ($N=35$ in V3 spin mode). The synodic frequency relative to the Earth is therefore the orbital plus the spin frequency, i.e. the frequency of the signal ($\nu_{{}_{EP}}=3.11133\,\rm mHz$ in V3 mode).
@@
An analysis of large glitches (with $\rm SNR>3$) carried out for a typical measurement session in Ref. MicroscopeGlitchesCQG2022-008 shows the FFT of these glitches in Fig. 7 which is worth reporting here (see Fig. 2). It shows that the large glitches occur at all the frequencies listed in Table 3 at the level of $10^{-11}\,\rm ms^{-2}$ and above, reaching about $7\times 10^{-11}\,\rm ms^{-2}$ at the spin frequency, i.e. the synodic frequency relative to the Sun. Higher harmonics appear too, but they are less relevant to the experiment. The accelerations occur simultaneously on the test cylinders of SUEP and SUREF sensor units, but they are not the same on the two TMs of each unit (the force is non gravitational and the masses are different), therefore the largest effect dominates the acceleration difference.
Figure 2: Same as Fig. 7 in MicroscopeGlitchesCQG2022-008 , where the caption reads: “Fourier transform of the time distribution of $\rm SNR>3$ glitches for a typical measurement session.” Note that in the inset the frequencies labeled as the “spin” frequency and the “EP” signal frequency are the synodic frequencies relative to the Sun and the Earth, whose higher harmonics appear in the full plot (not labeled), along with the orbital frequency which is the difference of the two.
Fig. 2 shows that in logarithmic scale each line appears as a double line because the two synodic frequencies are close to each other (see Table 3), except the line at the orbital frequency which is the difference of the two. In the inset showing a zoom in the region of the spin frequency the scale is linear and the two synodic frequencies clearly appear as distinct lines.
In Ref. MicroscopeGlitchesCQG2022-008 it is concluded that glitches must be removed from the data because it is impossible to accurately model their effects which include one at the frequency of the signal. In the time series of each session glitches are identified as all outliers above $4.5\,\sigma$ from the moving average, some time is allowed after each outlier to account for its decay, and all data points identified in this way are removed, up to 35% in SUEP and up to 40% in SUREF. As a result, the time series of the remaining data is interspersed with gaps. Gaps are then filled with artificial reconstructed data so that the resulting time series can be analyzed in the frequency domain to compute the spectral density of each measurement session, hence the limit to WEP violation $\delta_{i}\pm\sigma_{i}$. By including all $19$ SUEP sessions the final limit $\eta=\delta_{fin}\pm\sigma_{fin}=(-1.5\pm 2.3)\times 10^{-15}$-PRL2022 is obtained, where TouboulCQG2022-9 :
@@
The first one is related to the fact that gaps are distributed in correspondence of the outliers, which occur at the frequencies of Table 3, including the frequency of the signal; hence gaps too occur at these frequencies. Thus, gaps retain memory of the glitches whose elimination originated them; and since the sign of WEP violation is not known, they may in fact mimic a violation signal or cancel an effect (signal or systematic).
The second issue is that the artificial data at the level of random platform noise which replace the outliers do contribute to increasing the integration time of each measurement session, hence to reducing the level $\delta_{i}$ of WEP violation of the session. Then, according to (9), the final result is also reduced. If this procedure were correct, the more outliers are present, the larger the fraction of gaps, the larger the fraction of artificial data, the better the WEP test of the session and the WEP test of the entire experiment. Which is clearly a paradox. No matter how artificial data are generated and/or manipulated, they should not introduce any physical information, i.e. their effect on the $\delta$ estimated in each session ought to be null.
An interesting comparison is reported in MicroscopeGlitchesCQG2022-008 between the FFT of the differential acceleration obtained from the original measured data, and the FFT after glitches were removed and the missing data were reconstructed. This was done for SUREF session #380 lasting 120 orbits. This session has been chosen because, with equal composition test masses, any effect at $\nu_{{}_{EP}}$ is spurious, and the low noise makes the effects of glitches well visible.
Figure 3: Same as Fig. 10 in MicroscopeGlitchesCQG2022-008 , where the caption reads: “Spectra of SUREF’s $x$ axis differential acceleration for session 380, before (black) and after (red) glitches masking and data reconstruction.” ($x$ is the sensitive axis of the sensor)
Fig. 3, copied from Fig. 10 in MicroscopeGlitchesCQG2022-008 , shows the two plots, measured data in black, reconstructed data in red. All the double lines listed in Table 3 (plus the single line at $\nu_{{}_{orb}}$) are clearly visible (the tidal line at $2\nu_{{}_{EP}}$ has not been removed). It is apparent that the procedure used to eliminate the glitches only reduces their effects, including the one at $\nu_{{}_{EP}}$. In fact, in the final SUREF analysis this session, originally with $46\,\%$ of artificial data, has been split into two short sessions (of $46$ and $34$ orbits), each one with a small percentage of glitches ($7\,\%$ and $5\,\%$). Thus, most of the reconstructed data in Fig. 3 (red curve) have been eliminated, which is a reasonable decision.
@@
Yet, in all cases $\delta_{i}$ has been obtained with the same procedure, eliminating the same outliers, generating the same gaps, filling the gaps with the same data, the difference being only the fake signal which is therefore obviously recovered. The problem is that $\delta_{i}$ itsef may be affected by the artificial data, and this is not addressed by the check.
Instead, the only session with zero glitches (SUREF session #778-1) might be used, despite its short duration ($38$ orbits only) and large $\delta$ ($8.1\times 10^{-15}$ouboulCQG2022-9 , Table 6), for a different check: by inserting in the original time series a “fake” percentage of glitches with a distribution resembling the observed ones, then applying the elimination and data reconstruction procedure to recover the new $\delta$ for comparison with the original one. If successful, the check would give some confidence in the procedure.
@@
A clear-cut way out is to proceed in the footsteps of the Eöt-Wash ground tests of the WEP with rotating torsion balances Adel2008 ; AdelCQGfocusissue2012 . [^6] The amount of data removed due to sporadic spikes in the ion pump current, or due to abrupt deviations of the balance axis from the local vertical, was only 7% of the total Adel2008 . They established the limit to WEP violation using only the real measured data by exploiting the fact that the frequency and phase of a differential acceleration due to WEP violation in the field of a source body (the Earth in this case) are known, only its amplitude and sign are unknown.
@@
In the case of MICROSCOPE it works as follows. At any given time, in the reference frame rotating with the spacecraft, the position of the Earth and the phase of the sensitive axis are known. Hence, in the time series of the acceleration differences the violation signal is a sinusoid with the synodic frequency of the spacecraft relative to the Earth whose maximum amplitude (unknown) occurs twice per synodic period, when the sensitive axis points towards or away from Earth (sign unknown). This demodulated phase lock-in signal can be fitted to the time series of the acceleration differences –only the real measured ones– in order to determine its amplitude and sign, which therefore would not be affected by whatever gaps.
@@
A problem is that outliers have an effect at the frequency of the signal. However, we do not expect it to have the same phase. The differential acceleration of a potential WEP violation in the field of the Earth is a vector pointing to the center of mass of the Earth, hence it is maximum (in amplitude) when the coaxial cylinders of the sensor unit are along the radial direction. We have seen that glitches induced by heat from the Earth produce an acceleration difference at the same frequency, but they are certainly unrelated to the phase of the symmetry/sensitive axes (when the accumulated stress energy is released the test cylinders may be oriented in any direction relative to the Earth). Thus, should a residual spurious effect remain after the elimination of glitches, it would not be confused with the signal.
@@
This analysis applied to MICROSCOPE real measured data after removal of the outliers, would establish beyond question the result of the experiment.
V Conclusions
@@
A space test of the weak equivalence principle in the field of the Earth has enormous potential for a leap forward in precision, by building on a strong driving signal and better instrument isolation from nearby masses in space than in any ground laboratory, hence easier control of systematic effects from local disturbances PerBraginskyPLA2018 .
@@
A satellite test of the WEP has been over 40 years in the making, a number of designs were proposed, one -MICROSCOPE- has made it to flight, and has reported two orders of magnitude improvement over the best laboratory experiments to date MS-PRL2022 .
We summarize the salient features of our investigation of MICROSCOPE as follows.
i) The experiment configuration (two concentric coaxial cylinders, each of which is an independent accelerometer; capacitive position measurement and control; electric polarization and grounding by a thin gold wire; the displacements of each TM are measured from an origin –a capacitance zero– that must be newly determined on each measurement session; the corresponding accelerations are measured via the control force required to keep the TM at the capacitive zero; the acceleration difference between the two test masses of each sensor unit is computed after scale matching) gives rise to large accelerations biases, to be subtracted in post-processing, and a potentially unstable zero.
@@
ii) The experiment appears to be limited by thermal noise from internal damping in the grounding wires whose stiffness was predicted incorrectly in contrast with previous results from GOCE. Once in orbit, in order to reduce the noise level, the satellite spin was raised to a considerably larger rate than planned. The functional form of such noise is well known, and, lacking complete explicit information, we use the published data to calculate each session’s spectral density figure at the signal frequency, and express it in terms of the physical quantities it depends upon. We find jumps in sequential sessions; the corresponding quality factors vary widely and are sometimes unrealistically high or low. We suggest an unstable zero may be the cause of such jumps, an issue that would require careful investigation even if the grounding wire were to be replaced by an active discharger.
@@
iii) We find that a sub-$\mu\rm V$ voltage fluctuation, coupled to the large bias voltage made necessary by the capacitive control scheme, may cause the detected acceleration to leap in the way revealed by the data, and we suggest a potential cause in erratic patch charge rearrangements.
@@
iv) The experiment was affected by a large number of glitches, anomalous releases of energy, originating in the spacecraft, producing large acceleration differences at the signal and spin frequencies, their harmonics and their difference. We argue the form of such glitches supports the hypothesis they were triggered by cycles of heating and cooling, by the Sun and Earth, of the loose insulation foils. We find evidence that the way the glitches were treated –removed and replaced with artificial reconstructed data– leaves extant effects at the critical frequencies, including the frequency of the signal, that could mimic a violation signal or cancel an effect (systematic error or real signal). We argue that an analysis as successfully carried out in the Eöt-Wash laboratory tests would be tolerant of the missing data without affecting the sensitivity of the experiment.
@@
Any future WEP experiment in space must prove by its very design that it will not suffer from glitches (GOCE showed the way), and must avoid introducing artificial data potentially compromising the analysis. Over and above that, future experiments aiming to exploit the full potential of space must readdress the experiment design: null experiment, stable zero, minimal thermal noise by fast spin MicroscopeLimitations ; PerBraginskyPLA2018 .

### v4

@@ IV Limiting noise, patch charges and unstable zero
Using Eqs.(7) and (8), it leads to the relationship
@@
As recalled in Sec. III, the difference of acceleration biases between the TMs of SUEP is $\simeq 10^{-6}\,\rm ms^{-2}$. According to Eq.(10), a patch fluctuating potential $\delta V_{fpp}\simeq 0.3\,\mu V$ at the V3 signal frequency $\nu_{{}_{EP}}=3.11\,\rm mHz$ would produce a spurious acceleration of $4\times 10^{-14}\,\rm ms^{-2}$, as large as the average acceleration difference measured in all SUEP sessions (once normalized by session duration) at $\nu_{{}_{EP}}$ (Ref.(15), Fig. 10) and enough to change it by $100\%$ either way.
In GOCE a source of low frequency acceleration noise with $1/\sqrt{\nu}$ dependence below $0.01\,\rm Hz$ was found to be due to fluctuating potentials generated by charge patches on the test mass and its electrodes.14 GOCE was set up as a gravity gradiometer with $0.5\,\rm m$ separation of the TMs and a measurement bandwidth from $500\,\rm mHz$ to $0.1\,\rm Hz$. At $3\,\rm mHz$ the acceleration difference sensitivity was of $2\times 10^{-12}\,\rm ms^{-2}/\sqrt{Hz}$ along the ultra sensitive $Y,Z$ axes (more than one order of magnitude better than that of SUEP sensitive axis) and $4\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ along the least sensitive $X$ axis, with a contribution from patch fluctuating potentials of about $4\times 10^{-13}\,\rm ms^{-2}/\sqrt{Hz}$ and $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ respectively.25,  [^4] With no wire the zero of each TM control (dictated by geometry, construction, mounting and electronics) does not coincide with the zero of other forces acting on the test mass, such as Earth tides and electrostatic stiffness. Applying a constant bias is therefore inevitable in the Microscope design, regardless of the gold wire. Hence, any fluctuation of the bias in the frequency region of the WEP violation signal is relevant when aiming at very high precision.
@@
An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments, including those with laser tracked satellites, celestial bodies and cold atoms 26; 27; 28. In other tests, primarily those based on torsion balances, the physical observable is a true null.29
V Outliers, gaps and artificial data
@@
Since 2001 five space missions carrying ONERA’s accelerometers have been launched with a total of seven satellites, and all but GOCE have experienced such spikes. There is wide consensus that the spikes are triggered by energy inputs from the Earth and the Sun causing micro-vibration, e.g. release of stress energy in the spacecraft multi-layer insulation (MLI). At the time of the design of GOCE, reports of such effects caused alarm, and countermeasures were adopted at design and test level based on a stiffer insulation, which were successful, and no spikes were seen.
@@
In Microscope the transfer function of the two test cylinders in the sensor unit are not the same; in addition, the force is non gravitational and the masses are different. As a result, spikes are observed in the acceleration differences of each SU.
Microscope’s synodic frequency relative to the Sun is well approximated by the spin frequency relative to the fixed stars, the difference being the annual orbital frequency of the Earth around the Sun of $2\times 10^{-7}\,\rm rad/s$. The synodic frequency relative to the Earth is the orbital plus the spin frequency, i.e. the frequency of the signal ($\nu_{{}_{EP}}=3.11133\,\rm mHz$ in V3 mode).
@@
An FFT analysis of large glitches (with $\rm SNR>3$) for a typical measurement session is reported in Fig. 7 of Ref. (30). It shows that the large glitches occur at all the frequencies listed in Table 3 at the level of $10^{-11}\,\rm ms^{-2}$ and above, reaching about $7\times 10^{-11}\,\rm ms^{-2}$ at the spin frequency, i.e. the synodic frequency relative to the Sun. Higher harmonics appear too, but they are less relevant to the experiment. In the logarithmic scale of the Figure each line appears as a double line because the two synodic frequencies are close to each other (see Table 3), except the line at the orbital frequency which is the difference of the two. However, in the inset of the same Figure showing a zoom in the region of the spin frequency, the scale is linear and the two synodic frequencies clearly appear as distinct lines.
In Ref.(30) it is concluded that glitches must be removed from the data because it is impossible to accurately model their effects which include one at the frequency of the signal. In the time series of each session glitches are identified as all outliers above $4.5\,\sigma$ from the moving average, some time is allowed after each outlier to account for its decay, and all data points identified in this way are removed, up to 35% in SUEP and up to 40% in SUREF. As a result, the time series of the remaining data is interspersed with gaps. Gaps are then filled with artificial reconstructed data so that the resulting time series can be analyzed in the frequency domain to compute the spectral density of each measurement session, hence the limit to WEP violation $\delta_{i}\pm\sigma_{i}$. By including all $19$ SUEP sessions the final limit $\eta=\delta_{fin}\pm\sigma_{fin}=(-1.5\pm 2.3)\times 10^{-15}$ is obtained (Ref.(4)), where:8
@@
The first one is related to the fact that gaps retain memory of the glitches whose elimination originated them; and since the sign of WEP violation is not known, they may in fact mimic a violation signal or cancel an effect, signal or systematic.
The second issue is that the artificial data at the level of random platform noise which replace the outliers do contribute to increasing the integration time of each measurement session, hence to reducing the level $\delta_{i}$ of WEP violation of the session (Eq.6). Then, according to Eq.(11), the final result is also reduced. If this procedure were correct, the more outliers are present, the larger the fraction of gaps, the larger the fraction of artificial data, the better the WEP test of the session and the WEP test of the entire experiment. Which is clearly a paradox. No matter how artificial data are generated and/or manipulated, they should not introduce any physical information, i.e. their effect on the $\delta$ estimated in each session ought to be null.
An interesting comparison is reported in Ref.(30) between the FFT of the differential acceleration obtained from the original measured data, and the FFT after glitches were removed and the missing data were reconstructed. This was done for SUREF session #380 lasting 120 orbits. This session has been chosen because, with equal composition test masses, any effect at $\nu_{{}_{EP}}$ is spurious, and the low noise makes the effects of glitches well visible.
Fig. 10 of Ref.(30), shows the two plots, FFT of the measured data in black, of reconstructed data in red. All the double lines listed in Table 3 (plus the single line at $\nu_{{}_{orb}}$) are clearly visible (the tidal line at $2\nu_{{}_{EP}}$ has not been removed). It is apparent that the procedure used to eliminate the glitches only reduces their effects, including the one at $\nu_{{}_{EP}}$. In fact, in the final SUREF analysis this session, originally with $46\,\%$ of artificial data, has been split into two short sessions (of $46$ and $34$ orbits), each one with a small percentage of glitches ($7\,\%$ and $5\,\%$). Thus, most of the reconstructed data have been eliminated, which is a reasonable decision.
@@
Yet, in all cases $\delta_{i}$ has been obtained with the same procedure, eliminating the same outliers, generating the same gaps, filling the gaps with the same data, the difference being only the fake signal which is therefore obviously recovered. The problem is that $\delta_{i}$ itsef may be affected by the artificial data, and this fact is not addressed by the check.
@@
A clear-cut way out is to proceed in the footsteps of the Eöt-Wash ground tests of the WEP with rotating torsion balances. They established the limit to WEP violation using only the real measured data by exploiting the fact that the frequency and phase of a differential acceleration due to WEP violation in the field of a source body (the Earth in this case) are known, only its amplitude and sign are unknown.
@@
In the case of Microscope it works as follows. At any given time, in the reference frame rotating with the spacecraft, the position of the Earth and the phase of the sensitive axis are known. Hence, in the time series of the acceleration differences the violation signal is a sinusoid with the synodic frequency of the spacecraft relative to the Earth whose maximum size (of opposite sign and unknown modulus) occurs twice per synodic period, when the sensitive axis points towards or away from Earth (sign unknown). This demodulated phase lock-in signal can be fitted to the time series of the acceleration differences –only the real measured ones– in order to determine its amplitude and sign, which therefore would not be affected by whatever gaps.
@@
A problem is that in Microscope outliers have an effect at the frequency of the signal. However, we do not expect it to have the same phase. The differential acceleration of a potential WEP violation in the field of the Earth is a vector pointing to the center of mass of the Earth, hence it is maximum (in amplitude) when the coaxial cylinders of the sensor unit are along the radial direction. We have seen that glitches induced by heat from the Earth produce an acceleration difference at the same frequency, but they are certainly unrelated to the phase of the symmetry/sensitive axes (when the accumulated stress energy is released the test cylinders may be oriented in any direction relative to the Earth). Thus, should a residual spurious effect remain after the elimination of glitches, it would not be confused with the signal.
@@
This analysis applied to Microscope real measured data after removal of the outliers would establish beyond question the result of the experiment.
VI Conclusions
@@
A space test of the weak equivalence principle in the field of the Earth has enormous potential for a leap forward in precision, by building on a strong driving signal and better instrument isolation from nearby masses in space than in any ground laboratory, hence easier control of systematic effects from local disturbances.29
@@
A satellite test of the WEP has been over 40 years in the making, a number of designs were proposed, one –Microscope– has made it to flight, and has reported about two orders of magnitude improvement over the best laboratory experiments to date.
@@
The experiment appears to be limited by thermal noise from internal damping in the grounding wires whose stiffness was predicted incorrectly in contrast with previous results from GOCE. Once in orbit, in order to reduce the higher level of noise, the spin rate of the satellite was raised to a considerably larger rate than planned, only to find an unexplained noise jump. We find jumps, up or down, also in sequential sessions at the same spin rate, which are inexplicable on the basis of the current knowledge of internal damping noise and the physical quantities it depends upon.
@@
We find that voltage fluctuations at sub-$\mu\rm V$ level, coupled to the large bias voltage made necessary by the capacitive control scheme, may cause the detected acceleration to leap in the way revealed by the data, and we suggest a potential cause in erratic patch charge rearrangements leading to an unstable zero. An issue to be taken into account regardless of the gold wire.
@@
The experiment was affected by a large number of glitches, anomalous releases of energy, originating in the spacecraft, producing large acceleration differences at the signal and spin frequencies, their harmonics and their difference. We argue the form of such glitches supports the hypothesis they were triggered by cycles of heating and cooling, by the Sun and Earth, of the loose insulation foils. We find evidence that the way the glitches were treated –removed and replaced with artificial reconstructed data– leaves extant effects at the critical frequencies, including the frequency of the signal, that could mimic a violation signal or cancel an effect (systematic error or real signal). We argue that an analysis as successfully carried out in the Eöt-Wash laboratory tests would be tolerant of the missing data without affecting the sensitivity of the experiment.
@@
Any future WEP experiment in space must prove by its very design that it will not suffer from glitches (GOCE showed the way), and must avoid introducing artificial data potentially compromising the analysis.
Our investigation, having identified the strong and weak points of the first space test of the WEP, shows the way to the desirable properties of an experiment design targeting yet another leap in precision, to $10^{-17}$ or better: differential instrument; well-defined and stable zero point; large gaps between test masses; rotation about a symmetry axis; 2D sensitivity in the plane of the signal; fast spin.